OpenStudy (anonymous):

Jonny is able to invest $600 per month in an ordinary annuity (retirement account) at 7% compounded monhtly. After 10 years, the economy tanks. For the next 10 years, he can only afford $350 per month and the interest rates drop to 5% compoundly monthly during the second 10 years. How much total money does Johny have in his retirement account after 20 years?

OpenStudy (anonymous):

I would say that this is the result of 2 FV calculations. 1. Find the FV for the first 10yr period (use excel's FV function and input the following values: Rate = .07/12 = .00583333 Nper = 10*12 = 120 Pmt = -600 Resulting in a FV of 103,850.88 2. Preform the same calculation but change the Pmt amount to -350 and include a -103,850.88 for PV so: Rate = .07/12 = .00583333 Nper = 10*12 = 120 Pmt = -350 PV = -103,850.88 This should give you an overall FV of $269,284.79

OpenStudy (matthewrlee):

7% is great! Who is Johnny using?

OpenStudy (anonymous):

the first investment still collects interest during the second ten year period just at a 5% rate instead if 7%. So wouldn't it be[ 103,850.88( 1+ .05/12)^120]. Then add that amount to the second amount?

OpenStudy (anonymous):

Yes, thats why you include the -103,850.88 in as the PV of the investment. Because that is what you are starting out with and interest will accrue on that portion also.